OpenStudy (anonymous):

How to solve qubic equation: r^3 +3r^2 + 3r + 1 = 0 The calculator shows answer to be -1, but i need to know the steps, please solve with steps.

OpenStudy (anikate):

ok

OpenStudy (anikate):

r(r^2+3r+3)+1=o

OpenStudy (anikate):

do u know the x method

OpenStudy (anonymous):

co the durand-kerner method

OpenStudy (anonymous):

this problem is difficult

OpenStudy (precal):

can't you just use synthetic division to show that -1 is a zero?

OpenStudy (precal):

that is what I would do

OpenStudy (precal):

also then I am left with a quadratic which is easily solvable.

OpenStudy (anonymous):

i think you can write as \[\Large (r+1)^3=0\]

OpenStudy (anonymous):

which means \[\Large (r+1)(r+1)(r+1)=0\]

OpenStudy (anonymous):

now \[\Large r+1=0\] \[\Large r+1=0\] \[\Large r+1=0\] solving \[\Large r=-1\] so you have three same roots \[\Large r=-1\] \[\Large r=-1\] \[\Large r=-1\]

OpenStudy (precal):

sami-21 is correct

OpenStudy (precal):

my approach would also work

OpenStudy (anonymous):

note i used the following \[\Large (a+b)^3=a^3+b^3+3a^2b+3ab^2\] to write \[\Large (r^3+3r^2+3r+1)=(r+1)^3\] i hope it is clear

OpenStudy (anonymous):

@onaogh did you get it

OpenStudy (anonymous):

@sami-21, suppose we didn't have calculators, and you didn't knew that the roots were -1, how would u solve it then ??

OpenStudy (anonymous):

@onaogh i did solve it without any calculator !

OpenStudy (anonymous):

even if you want to know then the one factor must be factor of the constant term here the constant term is +1 its factor can be +1 or -1 so use both of them in them equation if by any value it gets zero it means it is root.

OpenStudy (precal):

use the p/q possible rationals thm that would lead you to +1 or -1

OpenStudy (anonymous):

can you try this: solve this: r^3 +19r^2 + 6r + 7 = 0

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

ok in this case old methods wouldn't work because of imaginary roots .

OpenStudy (anonymous):

i hope this was not in your book ??

OpenStudy (precal):

I would test 7, -7, 1, -1 as possible factors

OpenStudy (anonymous):

Suppose I have 4 unique integer roots, what is the strategy to guess the roots, I thought Vieta and Rational Roots would help?

OpenStudy (anonymous):

like what i said do duran-kerner method

OpenStudy (anonymous):

no, i made up that equation, i want to see how to be able to solve with any coefficients. @CJVelasco, yeah googling around and found few methods to solve.

OpenStudy (anonymous):

the answers would be in decimal

OpenStudy (precal):

for 4 your equation has to be a fourth power

OpenStudy (anonymous):

i knew that :) thats why i asked it cannot be from book :)

OpenStudy (anonymous):

thanks guys for your time. @sami-21 , thanks a lot

OpenStudy (precal):

btw imaginary numbers are valid solutions aka as complex solutions traditional methods would reveal them

OpenStudy (anonymous):

thanks @precal you are right ! traditional methods works for cube roots ! and can be found easily .but in the question given above it was not easy .i wrote the sentence in the context of the given question. Newton method is very fast and good method to approximate the rots of higher order polynomials.!